# Not all Layers of LLMs are Necessary during Inference

### v1

@@ Abstract
Siqi Fan2, Xin Jiang1, Xiang Li1, Xuying Meng3, Peng Han2, Shuo Shang2*, Aixin Sun4, Yequan Wang1*, Zhongyuan Wang1 ${}^{1}$eijing Academy of Artificial Intelligence, Beijing, China ${}^{2}$niversity of Electronic Science and Technology of China, Chengdu, China ${}^{3}$nstitute of Computing Technology, Chinese Academy of Sciences, Beijing, China ${}^{4}$chool of Computer Science and Engineering, Nanyang Technological University, Singapore
@@ 1 Introduction
To the best of our knowledge, this is the first attempt to design an early stop strategy for LLM inference. Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer can save an average of 14.8% computational resources, even up to 50% on sentiment tasks while maintaining comparable performance. More importantly, AdaInfer is orthogonal to other model acceleration techniques, offering the potential for further enhancing inference efficiency (Section 4).
@@ 3 AdaInfer
The workflow of AdaInfer and the computational efficiencies gained through this method are depicted in Figures 1(a) and 1(b), respectively. The key of AdaInfer is how to find the early stop signal while keeping the original abilities of LLMs. AdaInfer dynamically computes the stopping signal by evaluating critical features (i.e., “gap” and “top prob”). This process involves two main components: a Feature Selection module and a Classifier module. At each layer, the Feature Selection crafts a feature vector for the current input instance. Subsequently, the Classifier (often SVM or CRF for their effectiveness) assesses the stopping signal’s strength. A strong enough signal triggers an early process termination, allowing for the bypass of subsequent decoder layers.
@@ 3.1 Feature Selection
These analyses are showcased in Figure 3. In this figure, we observe the following trends: (1) For Llama2-13B with 40 layers Touvron et al. (2023) across sentiment and MMLU tasks, the “gap” and “top prob” gradually increase during the inference phase, stabilizing in the deeper layers. (2) The activation of “gap” and “top prob” varies across layers for different tasks. These phenomenons are also evident in the Llama2-7B, OPT-13B Zhang et al. (2022), and GPT-J Wang and Komatsuzaki (2021) (See Appendix C). This demonstrates gap and top prob can serve as universal features, indicating the stopping signal. Notably, “gap” and “top prob” values remain consistent across diverse tasks, suggesting a versatile classifier applicable to various tasks. Factor study in subsequent experiments also shows that other features exhibit subtle differences across layers.
@@ 3.2 Classifier
The Classifier determines if the signal is compelling enough to warrant an early termination of the process. Using confidence-based criteria doesn’t require extra computations during inference but may involve threshold tuning with validation data to balance accuracy and efficiency Huang et al. (2017); Yang et al. (2020). This involves a trade-off between performance and computational efficiency. Conversely, the plug-and-play nature of the gating function Lin et al. (2017); Bejnordi et al. (2019) provides greater universality. Nonetheless, discrete decision functions, lacking gradient information, often require specialized training methods.
@@ 4.1 Evaluation Tasks
GPT-3’s Brown et al. (2020) few-shot learning capability is tested with tasks requiring pattern recognition, using synthetic datasets from Todd et al. (2024); Hernandez et al. (2024) for tasks like Capitalize/Lowercase Letter, Choose Item/Category from List, and recognizing data pairs (e.g.,, Person-Occupation).
@@ 4.3 Comparison with Baseline Methods
Tables 2 and 3 show that across both zero-shot and few-shot settings, top-1 average accuracy remains within a narrow margin of $<$5% for all tasks and $<$1% for QA and text classification task groups when compared to baseline models. AdaInfer maintains mainstream LLM capabilities and in-context learning abilities without modifying model parameters. This finding is promising, especially in light of our observation1 in Section 2.2, where we demonstrate the feasibility of implementing early exit strategies within LLM middle layers while preserving accuracy. For certain tasks, AdaInfer surpasses the last layer (baseline) in zero-shot or few-shot accuracy. This hints at a tendency for deep layers to potentially over-represent certain tasks, which could impede performance during LLM inference.
Reducing FLOPs Costs from 2% to 41%.
@@
Task Setting AdaInfer w. GAP AdaInfer w. CRF
@@ 4.4 Evaluation on Different Exit Strategy
In the main experiments Table 2 and Table 3, we employed SVM as the classifier for AdaInfer. To explore the impact of different classification strategies, Table 4 compares the effects of implementing an early-exit strategy with a GAP threshold set at 0.8 (stopping computation when the current block’s GAP feature exceeds 0.8) against using CRF as a classifier. The results indicate that both GAP and CRF can reduce computational costs from 3% to 50% and maintain comparable LLM performance. Notably, in the zero-shot setting, GAP outperforms CRF, suggesting a relatively weak dependency between block features.
@@ 4.5 Evaluation across Scaling Law
In our main experiments, we employed 13B-sized Llama2 and OPT models. To explore the effects of AdaInfer on models of different sizes, we conducted experiments on the 7B and 70B versions of Llama. The results for the 7B model, presented in Table 5, show that AdaInfer either maintains accuracy with minimal (<1%) loss or exceeds the baseline in certain tasks, and achieves a computational reduction ranging from 4% to 32%. However, in experiments with the 70B model, we observed that in a zero-shot setting, AdaInfer matches or slightly exceeds the baseline model while reducing computational costs by 10% to 50%. However, in the few-shot setting, despite similar reductions in computation, AdaInfer’s accuracy shows a 1% to 25% drop across different tasks compared to the baseline. This suggests that for larger models, such as the 70B or even larger scales, AdaInfer may need to more precisely identify and utilize features at different levels. Improving AdaInfer to adapt to these larger models is a direction for our future research. The results of all LLMs using different classifiers are summarized in Table 8 and Table 9 in the Appendix D and we have highlighted the best results for each task in the current setting.
@@ 4.6 Generalization Study
In Table 2 and Table 3, we randomly selected 6 to 9 training datasets from all task training sets (note: there are 71 subdatasets in total). Furthermore, to assess the generalization performance of the statistical classifiers, we conducted the following tests.
@@ 4.7 Factor Study
In response to the features mentioned in Section 3.1, we conducted cross-validation. Given that the classifiers in the main results utilized basic features (i.e., “gap”, “top prob”), we explored the impact of features such as the cosine similarity between the current block and the previous block, which encompasses the attention values (attn), multi-layer perceptron (mlp), and hidden states. The results are presented in Table 7. The attention values have no discernible impact on the results, while other features like mlp and hidden states have an adverse effect. This result is consistent with the trend shown in Figure 3, indicating that logits can measure whether the model’s current forward progress is sufficient, while changes in other features may involve various factors.
@@
5 Related Work
A straightforward approach to address the observation outlined in Section 2 involves dynamic neural networks Han et al. (2021); Huang et al. (2017); Bolukbasi et al. (2017). Consequently, networks with dynamic structures can be classified into two types: dynamic depth (number of network layers) and width (number of channels, parallel subnetworks, etc.). Dynamic depth integrates classifiers into intermediate layers of CNN/DNN networks for visual tasks to handle early exits Bolukbasi et al. (2017); Huang et al. (2017); Teerapittayanon et al. (2016), our proposed AdaInfer closely aligns with this concept. Meanwhile, skip-layer functionality dynamically omits the execution of a layer (or module) for any input token, facilitated by a gate function Wang et al. (2018) or a binary router Zeng et al. (2023). Other model accelerate methods such as quantization Xiao et al. (2023); Frantar et al. (2022), sparsity Liu et al. (2023) or distillation Touvron et al. (2021) are orthogonal areas and usually excel in different settings.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China (2022ZD0116300) and the National Science Foundation of China (NSFC No. 62106249).
@@ Appendix D Comprehensive Summary of Results
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract
Siqi Fan2, Xin Jiang1, Xiang Li1, Xuying Meng3, Peng Han2, Shuo Shang2*, Aixin Sun4, Yequan Wang1*, Zhongyuan Wang1 1Beijing Academy of Artificial Intelligence, Beijing, China 2University of Electronic Science and Technology of China, Chengdu, China 3Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 4School of Computer Science and Engineering, Nanyang Technological University, Singapore
@@ 1 Introduction
Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer can save an average of 14.8% computational resources, even up to 50% on sentiment tasks while maintaining comparable performance. More importantly, AdaInfer is orthogonal to other model acceleration techniques, offering the potential for further enhancing inference efficiency (Section 5).
2 Related Work
A straightforward approach to achieve adaptive inference involves dynamic neural networks Han et al. (2021); Huang et al. (2017); Bolukbasi et al. (2017). Consequently, networks with dynamic structures can be classified into two types: dynamic depth (number of network layers) and width (number of channels, parallel subnetworks etc.).
Dynamic depth.
Dynamic depth involves two methods: Early Exit (EE) and Skip layer. EE first appeared in CNN/DNN networks for visual tasks Bolukbasi et al. (2017); Huang et al. (2017); Teerapittayanon et al. (2016). Subsequently, it was utilized in accelerating the inference of encoder-only architectures in BERT by Li et al. (2020); Liu et al. (2020); Li et al. (2021); Kong et al. (2022). Recently, Schuster et al. (2022); Varshney et al. (2023) discuss confidence-based EE for LM adaptive inference. Our proposed AdaInfer closely aligns with EE concept. Specifically, we apply EE to mainstream decoder-only LLMs, which adhere to the scaling law but suffer from high inference costs due to their large parameter count. Meanwhile, skip-layer dynamically omits the execution of middle layers (or modules) for any input token, facilitated by a gate function Wang et al. (2018) or a binary router Zeng et al. (2023) and layer pruning Kim et al. (2024); Yang et al. (2024); Song et al. (2024). The main difference between our method and theirs is that we achieve instance-wise inference without altering the model parameters, which is crucial for current LLMs. To the best of our knowledge, this is the first attempt to discover that each block’s logits are crucial elements for EE classifiers in LLMs, and we incorporate it as a fundamental design choice in AdaInfer.
Dynamic width.
Dynamic width controls the number of neurons on the network width for efficient inference, such as reducing the number of CNN channels Hua et al. (2019); Hoefler et al. (2021) and establishing multiple parallel structures for “experts” in MoE Fedus et al. (2022); Zhou et al. (2022); Artetxe et al. (2021), dynamically weighting and predicting the output results. Recently, Ma et al. (2023); Addanki et al. (2023); Xia et al. (2023) slimming the network width by pruning attention heads and the output neuron in Query, Key or Value. Other model accelerate methods such as quantization Xiao et al. (2023); Frantar et al. (2022), sparsity Liu et al. (2023) are orthogonal areas and usually excel in different settings.
@@ 4 AdaInfer
The workflow of AdaInfer and the computational efficiencies gained through this method are depicted in Figures 2(a) and 2(b), respectively. The key of AdaInfer is how to find the early stop signal while keeping the original abilities of LLMs. AdaInfer dynamically computes the stopping signal by evaluating critical features (i.e., “gap” and “top prob”). This process involves two main components: a Feature Selection module and a Classifier module. At each layer, the Feature Selection crafts a feature vector for the current input instance. Subsequently, the Classifier (often SVM or CRF for their effectiveness) assesses the stopping signal’s strength. A strong enough signal triggers an early process termination, allowing for the bypass of subsequent decoder layers.
@@ 4.1 Feature Selection
These analyses are showcased in Figure 3. In this figure, we observe the following trends: (1) For Llama2-13B with 40 layers Touvron et al. (2023) across sentiment and MMLU tasks, the “gap” and “top prob” gradually increase during the inference phase, stabilizing in the deeper layers. (2) The activation of “gap” and “top prob” varies across layers for different tasks. These phenomenons are also evident in the Llama2-7B, OPT-13B Zhang et al. (2022), and GPT-J Wang and Komatsuzaki (2021) (See Appendix C). This demonstrates “gap” and “top prob” can serve as universal features, indicating the stopping signal. Notably, these two values remain consistent across diverse tasks, suggesting a versatile classifier applicable to various tasks. Factor study in subsequent experiments also shows that other features (e.g., Cosine Similarity) exhibit subtle differences across layers.
@@ 4.2 Classifier
The Classifier determines if the signal is compelling enough to warrant an early termination of the process. The rule-based approach heavily relies on rules, and the cost of individually constructing domain-specific features is high  Huang et al. (2017); Yang et al. (2020); Wang et al. (2022). Conversely, the plug-and-play nature of the gating function Lin et al. (2017); Bejnordi et al. (2019) provides greater universality. Nonetheless, discrete decision functions, lacking gradient information, often require specialized training methods.
@@ 5.1 Evaluation Tasks
GPT-3’s Brown et al. (2020) few-shot learning capability is tested with tasks requiring pattern recognition, using synthetic datasets from Todd et al. (2024); Hernandez et al. (2024) for tasks like Capitalize/Lowercase Letter, Choose Item/Category from List, and recognizing data pairs (e.g.,, Landmark-Country).
@@ 5.3 Comparison with Baseline Methods
Tables 2 and 3 show that across both zero-shot and few-shot settings, top-1 average accuracy remains within a narrow margin of $<$5% for all tasks and $<$1% for QA and text classification task groups when compared to baseline models. AdaInfer maintains LLMs capabilities and in-context learning abilities without modifying model parameters. This finding is promising, especially in light of our observation1 in Section 3.2, where we demonstrate the feasibility of implementing early exit strategies within LLM middle layers while preserving performance. For certain tasks, AdaInfer surpasses the last layer (baseline) in zero-shot or few-shot accuracy. This hints at a tendency for deep layers to potentially over-represent certain tasks, which could impede performance during LLM inference.
Reducing FLOPs Cost from 2% to 41%.
@@
Task Setting AdaInfer w. Rule AdaInfer w. CRF
@@ 5.4 Evaluation on Different Exit Strategy
In the main experiments Table 2 and Table 3, we employ SVM as the classifier for AdaInfer. To explore the impact of different classification strategies, Table 4 compares the effects of implementing an early-exit strategy with a GAP threshold set at 0.8 (stopping computation when the current block’s GAP feature exceeds 0.8) against using CRF as a classifier. The results indicate that both GAP and CRF can reduce computational costs from 3% to 50% and maintain comparable LLM performance. Notably, in the zero-shot setting, GAP outperforms CRF, suggesting a relatively weak dependency between block features.
@@ 5.5 Evaluation across Scaling Law
In our main experiments, we employ 13B-sized Llama2 and OPT models. To explore the effects of AdaInfer on models of different sizes, we conduct experiments on the 7B and 70B versions of Llama2. The results for the 7B model, presented in Table 5, show that AdaInfer either maintains accuracy with minimal (<1%) loss or exceeds the baseline in certain tasks, and achieves a computational reduction ranging from 4% to 32%. However, in experiments with the 70B model, we observe that in a zero-shot setting, AdaInfer matches or slightly exceeds the baseline model while reducing computational costs by 10% to 50%. However, in the few-shot setting, despite similar reductions in computation, AdaInfer’s accuracy shows a 1% to 25% drop across different tasks compared to the baseline. This suggests that for larger models, such as the 70B or even larger scales, AdaInfer may need to more precisely identify and utilize features at different levels. Improving AdaInfer to adapt to these larger models is a direction for our future research. The results of all LLMs using different classifiers are summarized in Table 8 and Table 9 in the Appendix D and we have highlighted the best results for each task in the current setting.
@@ 5.6 Generalization Study
In Tables 2 and 3, we randomly select 6 to 9 training datasets from the entire pool of task training sets, which altogether contain 71 sub-datasets, to train the AdaInfer classifier. Furthermore, to assess the generalization performance of the statistical classifiers, we conduct the following tests.
@@ 5.7 Factor Study
In response to the features mentioned in Section 4.1, we conduct cross-validation. Given that the classifiers in the main results utilized basic features (i.e., “gap”, “top prob”), we explore the impact of features such as the cosine similarity between the current block and the previous block, which encompasses the attention values (attn), multi-layer perceptron (mlp), and hidden states. The results are presented in Table 7. The attention values have no discernible impact on the results, while other features like mlp and hidden states have an adverse effect. This result is consistent with the trend shown in Figure 3, indicating that logits can measure whether the model’s current forward progress is sufficient, while changes in other features may involve various factors.
@@ Acknowledgments
This work is supported by the National Science and Technology Major Project (2022ZD0116300) and the National Science Foundation of China (No. 62106249).
@@ Appendix D Comprehensive Summary of Results
Zero-shot OPT-13B 7.95 100 8.20 100 20.00 100 12.05 100
AdaInfer w. Rule 3.21 89.58 0.60 85.17 20.72 87.98 8.18 87.58
AdaInfer w. CRF 7.14 96.57 4.60 93.26 24.36 93.22 12.03 94.35
AdaInfer 8.67 97.55 2.80 97.55 23.00 97.55 11.49 97.55
Few-shot OPT-13B 23.60 100 21.45 100 26.12 100 23.72 100
AdaInfer w. Rule 20.99 79.54 20.72 80.00 24.20 82.93 21.97 80.82
AdaInfer w. CRF 24.44 97.43 21.18 97.55 25.98 97.11 24.81 97.37
AdaInfer 22.59 83.94 21.62 86.05 25.95 88.31 23.39 86.10
Zero-shot Llama2-7B 4.19 100 5.30 100 20.40 100 9.96 100
AdaInfer w. Rule 4.69 95.69 4.60 94.90 23.90 89.48 11.06 93.36
AdaInfer w. CRF 4.86 95.32 2.00 95.01 18.80 91.17 8.55 93.83
AdaInfer 4.63 96.13 4.80 95.26 23.80 89.98 11.08 93.79
Few-shot Llama-2-7B 43.05 100 53.50 100 48.08 100 48.21 100
AdaInfer w. Rule 44.03 93.69 52.83 90.23 45.68 86.72 47.51 90.21
@@
AdaInfer w. Rule 5.35 90.84 1.10 92.78 24.60 73.17 10.35 85.60
AdaInfer w.CRF 4.77 97.40 1.40 97.28 23.10 93.03 9.76 95.90
AdaInfer 2.48 98.14 0.70 98.37 25.90 85.34 9.69 93.95
Few-shot Llama-2-13B 53.31 100 64.92 100 52.9 100 57.04 100
AdaInfer w. Rule 47.09 84.10 55.33 79.57 43.43 71.19 48.62 78.29
@@
AdaInfer 92.97 80.28 72.83 100 82.90 90.14 52.83 89.74
Zero-shot Llama2-7B 0.00 100 0.10 100 0.05 100 5.47 100
AdaInfer w. Rule 0.00 96.08 0.10 91.05 0.05 93.57 5.41 91.20
AdaInfer w. CRF 0.00 96.07 0.10 92.20 0.05 94.14 3.62 92.08